OpenStudy (anonymous):

Integrate :

OpenStudy (anonymous):

\[\huge \int\limits_{\frac{1}{\sqrt{2}}}^{\frac{\sqrt{3}}{2}} \ln e^{\cos x}d(\sin x)\]

OpenStudy (anonymous):

@ash2326 @Hero @Luigi0210 @Callisto @robtobey @dan815

OpenStudy (anonymous):

Substitute \(u=\cos x\), so \(du=-\sin x ~dx~\iff~-du=d(\sin x)\), if I'm interpreting the notation correctly.

OpenStudy (anonymous):

Or you can simply say \[\ln e^{\cos x}=\cos x\ln e=\cos x\]